Hello I have a word problem that i can't seem to figure out. I need a little help.
Taxpayers with an office in thir home may deduct a percentage of their home related expenses. This percentage is based on the ratio of the office's area to the area of the home. A taxpayer with an office in a 2200 square foot home maintains a 20 foot by 16 foot office. If the yearly electric bills for the home come to $4800, how much of this is deductible?
Thank you

Found 2 solutions by jorel1380, ccs2011:
Answer by jorel1380(3719)

The office is 16x20,so its area is 320 sq. ft.
In a 2200 sq. ft. home,that represents 320/2200 or 14.54% of the total area of the house. .14545454 x 4800=$698.18 of the electric bill.

Answer by ccs2011(207)

Area of office = 20*16 = 320
Ratio of office to home = 320/2200 = 8/55
As a percent 8/55 = 14.54%
Therefore 14.54% of his expenses are deductible
=> 4800*(.1454) = 698.18